OpenStudy (anonymous):

Write a formula for this sequence: 19, 11, 3, -5,... a. 27 - 8n b. -8n + 19 c. -5 + 8n

OpenStudy (anonymous):

B

OpenStudy (anonymous):

1st term A = 19, Common difference D = 11-19 = -8, So use nth term = A + (n-1)D .

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@DaWizjr nope. Infact it would be 19 + (n-1)-8 => 19-8n+8 which is 27 - 8n i.e. a

OpenStudy (anonymous):

Just typed A though